Consider the following transactions for Giggly Care.

Payment of dividends $1,250
Fees earned $4,500
Payment of expenses $2,750


Which of the following shows the effect of these transactions on the liquidity metric and the profitability metric of the company?

a.A decrease of $500 in the liquidity metric and an increase of $1,750 in the profitability metric

b.An increase of $500 in the liquidity metric and an increase of $1,750 in the profitability metric

c.An increase of $500 in the liquidity metric and a decrease of $1,750 in the profitability metric

d.An increase of $1,750 in the liquidity metric and a decrease of $500 in the profitability metric
Answer #1

Fees Earned =$4500

Payment of expense =$2750

Profit =4500-2750=$1750

Profitability metric is increased by $1750

Out of $1750 Profit , dividend paid =$1250

Increase in liquidity metric=1750-1250=$500

ANSWER:

b.An increase of $500 in the liquidity metric and an increase of $1,750 in the profitability metric